A study examined the association between lung cancer and working in an asbestos plant. Although the asbestos plant closed 10 years ago, researchers pulled past employment records to confirm employment. They divided the company employees into two groups: those directly exposed on the production line and those who worked in other areas of the plant, not directly working with asbestos. The researchers then looked at medical records to see if the exposed and unexposed groups were currently diagnosed with lung cancer. Of the 50 employees of the plant, 25 were exposed on the production line and 25 were not exposed. Based on current medical records, 19 of the exposed employees now have lung cancer and 8 of the unexposed now have lung cancer.

What is the study design?                             [ Select ]                       ["Ecologic", "Cross-sectional", "Case-control", "Prospective cohort", "Retrospective cohort", "Randomized control trial"]      

What is the answer to the main calculation of association in this study? (if it is a cohort or experimental study only produce the relative risk or risk ratio)                             [ Select ]                       ["Approximately 2.38", "Approximately 0.42", "Approximately 6.73", "Approximately 0.15", "None of the answers listed here are correct."]      

What is an important issue associated with this study design?                            [ Select ]                       ["Respondents may be lost to follow-up", "Temporal ambiguity", "It is hard to find cases to match controls", "The researcher may struggle with non-compliance", "None of the answers listed here are correct."]      
Answer #1

What is the study design?

Retrospective cohort

A retrospective study looks backwards and examines exposures to suspected risk or protection factors in relation to an outcome that is established at the start of the study

in this study we find the employers from the asbestose plant were find out and was assessing the development of lung cancer in them.

What is the answer to the main calculation of association in this study?

Approximately 2.38

Lung cancer

Yes

No

direct exposer to asbestos

Yes

19

6

No

8

17

RR = (A/(A+B))/(C/(C+D))

= (19/25)/(8/25)= 0.76/0.32

= 2.375

What is an important issue associated with this study design?

None of the answers listed here are correct